# spacing=nonfrench

### v1

@@ 8.4. Switch Deployment
NetCache (Jin et al., 2017) 8800 (57.3%) 12 (100%) 45 (93.8%) 528 (68.8%)
FarReach (Sheng et al., 2023) 8992 (58.5%) 12 (100%) 45 (93.8%) 499 (65.0%)
FMCache 8976 (58.4%) 12 (100%) 47 (97.6%) 712 (92.7%)
FMCache+ 8976 (58.4%) 12 (100%) 47 (97.6%) 712 (92.7%)
Table 3. (Exp#9) Switch resource usage (numbers in brackets refer to the fractions of used resources over total available ones). Numbers of NetCache and FarReach are quoted from (Sheng et al., 2023).
9. Related Work
@@
Scaling file-system metadata management. Metadata partitioning distributes file-system management across multiple servers for scalability. There are two primary approaches: (i) dynamic sub-tree partitioning, which distributes namespace sub-trees across servers (Weil et al., 2006; Sevilla et al., 2015; Wang et al., 2021a), and (ii) hash-based partitioning, which distributes file-system metadata across servers via hashing (Patil and Gibson, 2011; Thomson and Abadi, 2015; Niazi et al., 2017; Pan et al., 2021; Liao and Abadi, 2023; Wang et al., 2023). FMCache complements these approaches as an in-switch cache that absorbs operations upstream of file-system metadata layers.
@@
Client-side caching. PanFS (Welch et al., 2008) caches file and directory metadata and provides callbacks for cache consistency. IndexFS (Ren et al., 2014) and LocoFS (Li et al., 2017a) use lease-based caching for metadata management and invalidate cache entries upon lease expiration, but incur high overhead for renewing cache entries’ leases. InfiniFS (Lv et al., 2022) applies lazy invalidation for directory access metadata to limit the overhead of lease-based caching. Client-side caching often incurs high client-side complexity and overhead in maintaining cache consistency across a large number of clients. FMCache simplifies cache consistency by caching file-system metadata in a programmable switch that lies on the critical paths of multiple clients.
@@
In-switch caching. Programmable switches have been extensively studied for concurrency control (Li et al., 2017b; Zhu et al., 2019), network monitoring (Gupta et al., 2018), replication coordination (Jin et al., 2018), remote procedure calls (Zhao et al., 2023), key-value stream aggregation (He et al., 2023), and distributed lock management (Yu et al., 2020; Zhang et al., 2024). Several studies explore in-switch caching. SwitchKV (Li et al., 2016) and NetCache (Jin et al., 2017) design write-through in-switch caching for read-intensive workloads, and DistCache (Liu et al., 2019) designs distributed write-through caching across multiple switches. Pegasus (Li et al., 2020) and TurboKV (Eldakiky et al., 2021) cache replica-to-server mappings for replica selection. Mind (Lee et al., 2021) caches object-to-memory mappings for disaggregated memory systems. Concordia (Wang et al., 2021b) tracks cache copy locations and states to address concurrency in shared memory systems. FarReach (Sheng et al., 2023) designs fault-tolerant write-back caching for write-intensive workloads. The above in-switch caches are designed for key-value stores, which differ from file-system semantics. AsyncFS (Xu et al., 2024) proposes in-switch tracking of directory updates, while maintaining a client-side metadata cache for path resolution. In contrast, FMCache moves file-system metadata caching to switches, and also complements client-side caching.
10. Conclusion
@@
FMCache is an in-switch file-system metadata caching framework, aiming to achieve high throughput and load balancing for distributed file-system metadata services. It employs path-aware cache management, multi-level read-write locking, and local hash collision resolution. Experiments on a Tofino-switch testbed show that FMCache achieves significant throughput gains and complements client-side caching.

### v2

@@ VIII-D Switch Deployment
NetCache [jin17] 7856 (51.2%) 12 (100%) 45 (93.8%) 528 (68.8%)
FarReach [sheng25] 8080 (52.6%) 12 (100%) 45 (93.8%) 499 (65.0%)
Fletch 8976 (58.4%) 12 (100%) 47 (97.6%) 712 (92.7%)
Fletch+ 8976 (58.4%) 12 (100%) 47 (97.6%) 712 (92.7%)
IX Related Work
@@
Scaling file-system metadata management. Metadata partitioning distributes file-system management across multiple servers for scalability. There are two primary approaches: (i) dynamic sub-tree partitioning, which distributes namespace sub-trees across servers [weil06, sevilla15, wang21], and (ii) hash-based partitioning, which distributes file-system metadata across servers via hashing [patil11, thomson15, niazi17, pan21, liao23, wang23]. Fletch complements these approaches as an in-switch cache that absorbs operations upstream of file-system metadata layers.
@@
Client-side caching. PanFS [welch08] caches file and directory metadata and provides callbacks for cache consistency. IndexFS [ren14] and LocoFS [li17] use lease-based caching for metadata management and invalidate cache entries upon lease expiration, but incur high overhead for renewing cache entries’ leases. InfiniFS [lv22] applies lazy invalidation for directory access metadata to limit the overhead of lease-based caching. Client-side caching often incurs high client-side complexity and overhead in maintaining cache consistency across a large number of clients. Fletch simplifies cache consistency by caching file-system metadata in a programmable switch that lies on the critical paths of multiple clients.
@@
In-switch caching. Programmable switches have been extensively studied for concurrency control [li17eris, zhu19], network monitoring [gupta18], replication coordination [jin18], remote procedure calls [zhao23], key-value stream aggregation [he23], and distributed lock management [yu20, zhang24]. Several studies explore in-switch caching. SwitchKV [li16] and NetCache [jin17] design write-through in-switch caching for read-intensive workloads, and DistCache [liu19] designs distributed write-through caching across multiple switches. Pegasus [li20] and TurboKV [eldakiky20] cache replica-to-server mappings for replica selection. Mind [lee21] caches object-to-memory mappings for disaggregated memory systems. Concordia [wang21concordia] tracks cache copy locations and states to address concurrency in shared memory systems. FarReach [sheng25] designs fault-tolerant write-back caching for write-intensive workloads, and its extension, DistReach [sheng25], supports multiple switches. The above in-switch caches are designed for key-value stores, which differ from file-system semantics. SwitchFS [xu24] proposes in-switch tracking of directory updates, while maintaining a client-side metadata cache for path resolution. In contrast, Fletch moves file-system metadata caching to switches, and also complements client-side caching.
X Conclusion
@@
Fletch is an in-switch file-system metadata caching framework, aiming to achieve high throughput and load balancing for distributed file-system metadata services. It employs path-aware cache management, multi-level read-write locking, and local hash collision resolution. Experiments on a Tofino-switch testbed show that Fletch achieves significant throughput gains and complements client-side caching.